OpenStudy (sasogeek):

If 2tan(x)=1.5, find the value of cos(x)+2sin(x) If 2tan(x)=1.5, find the value of cos(x)+2sin(x) @Mathematics

OpenStudy (jamesj):

this means tan x = 3/4 and x is an angle either in the first or third quadrant. For the moment, let's assume it's on the first. Now sec^2 x = 1 + tan^2 x = 25/16 => cos x = 4/5 and hence sin x = 3/5. Thus cos x + 2 sin x = 2. Now find the answer if x is in the third quadrant.

OpenStudy (anonymous):

i didn't know

OpenStudy (sasogeek):

I got the same thing... just wanna be sure that I was right... thank you

OpenStudy (anonymous):

yup its right

OpenStudy (jamesj):

But be careful. +2 isn't the only answer because there's two locations of x.